OpenStudy (anonymous):

1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (anonymous):

T + B = 56 B = 2T +8 solve the system of equations

OpenStudy (anonymous):

T + B = 56 -2T + B = 8 subtract 3T = 48 T = 16 plug T back in 16 + B = 56 B = 40

OpenStudy (anonymous):

Belinda is 40 and Tony is 16

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

youre welcome